Two charge q &-2q are placed at some distnace the locus point in the place of the charge,where the potential is be zero?
A.-Circle
B.-Ellipse
3.-parabola

Grade:12

4 Answers

Sher Mohammad IIT Delhi
9 years ago
Ans is A:Circle
let charge q is at (0,0)and -2q at(r,0)
any point is (h,k)
potential is zero so kQq/sqrt(h^2+k^2)=2kQq/sqrt((h-r)^2+k^2)
=> 4(h^2+k^2)=((h-r)^2+k^2)
=>3h^2+3k^2+2rh-r^2=0 its circle

Sher Mohammad, IIT Delhi
faculty Askiitians

aswanth
9 years ago
Circle
 
aswanth
9 years ago
circle
Pinku
9 years ago
Answer is circle